A car travels 45 miles per hour in 20 minutes, how many miles has the car traveled?
Found 2 solutions by rfer, nyc_function:
Answer by rfer(16322)

45*1/3=45/3=15 miles

Answer by nyc_function(2741)

We first change 20 minutes to time in hours.

There are 60 minutes in 1 hour.

So, 20/60 = 1/3 hours.

We now use the formula D = rt.

We are searching for D = distance

D = (45)(1/3)

D = (15/1)

D = 15 miles

Done!